i got 2 questions. One motions one Newton’s Law

Q1) Object A is dropped from height h.At the same instant object B is thrown vertically upwards from the ground. Right before they collide in mid-air, the speed of A is twice of the speed of B. Determine the height h of collision

Q2) The pages of 2 identical books A and B are overlapping on one and another. The mass of each book is 1000g and the number of pages is \mu =0.3 . Book A is fixed on the table which a horizontal force F is pulling book B. Determine the minimum value of F to pull book B out.
1
Expert's answer
2020-03-19T11:38:16-0400

1)

For the object A

applying formula "v^2=u^2+2gS"

"v^2=0+2\times g\times h/2
\\v=\sqrt{gh}"


For the object B , speed at mid height

applying formula "v^2=u^2+2(-g)S"

"v^2=v_b^2-2\times g\times h/2\\v=\sqrt{v_b^2-gh}"


it is given that the speed of A is twice of the speed of B

"\sqrt{gh}=2\sqrt{v_b^2-gh}"

"gh=4(v_b^2-gh)\\5gh=4v_b^2\\h=\dfrac{4v_b^2}{5g}"

2) Normal force = mg=10N

maximum frictional force = "\mu"N =0.3x10=3N

so the minimum force needed to pull the book B is 3N